# A New Search Pipeline for Short Gamma Ray Bursts in Fermi/GBM Data - A 50% Increase in the Number of Detections

### v1

@@ Appendix I Contaminant Removal Tests Histograms
NameaaName: The trigger name as appears in the GBM trigger catalog. If it is a new trigger, the trigger time is displayed instead. T90 (s) SNR ($\sigma$)bbSNR ($\sigma$): The significance of the trigger at triggering, as appears in the GBM data product: TRIGDAT file. It is reported in units of standard deviations of standard normal distribution, hence it has the same meaning as our SNR. RA (deg) Dec (deg) Duration (s)ccDuration: The burst duration that had the maximal SNR in our search. SNR ($\sigma$) RA (deg)ddRA (deg): The maximum likelihood estimation of the right ascension, obtained by the sky posterior. Dec (deg)eeDEC (deg): The maximum likelihood estimation of the declination, obtained by the sky posterior. $E_{\text{peak}}$ (KeV)fffootnotemark: $\mathcal{B}_{\oplus}$ $\mathcal{B}_{\odot}$
bn140105065 1.09 6.40 208.22 50.17 0.242 21.44 210.17 49.79 366.86 0.00 0.00
bn140105748 0.58 4.70 252.88 19.03 3.606 19.15 256.83 23.13 495.30 0.00 0.00
bn140109771 0.70 4.80 102.74 29.76 0.327 16.21 195.73 41.64 668.72 0.00 0.00
bn140110411 0.77 4.60 50.64 -69.29 0.327 11.06 334.96 -76.88 366.86 0.30 0.20
bn140122597 3.58 4.80 56.08 15.09 2.671 17.58 58.98 3.16 110.41 0.00 0.00
2014-01-24 18:00:09.62 - - - - 0.133 11.90 222.52 -36.77 495.30 0.00 0.00
2014-01-29 07:06:50.56 - - - - 0.133 10.79 79.91 -6.35 201.26 0.00 0.00
bn140129499 0.13 8.20 183.40 -10.32 0.054 23.71 185.10 -3.97 149.06 0.00 0.00
bn140209313 1.41 5.40 81.33 32.49 0.596 235.61 89.66 23.69 110.41 0.00 0.00
2014-02-09 22:03:12.19 - - - - 2.671 9.78 281.68 -28.26 81.78 0.00 0.07
bn140211091 3.46 5.00 115.84 -13.59 1.979 18.46 115.91 -10.83 110.41 0.00 0.00
bn140216331 2.43 5.10 194.04 31.46 0.596 12.25 204.93 39.76 366.86 0.00 0.00
bn140224382 2.30 4.70 2.78 20.37 2.671 25.91 330.99 27.19 271.72 0.00 0.00
2014-02-25 04:03:20.64 - - - - 3.606 10.03 348.73 17.04 149.06 0.01 0.00
2014-03-03 06:44:35.49 - - - - 0.073 12.02 297.48 8.78 81.78 0.00 0.00
2014-03-07 00:43:25.41 - - - - 0.022 17.68 148.46 69.51 1645.78 0.09 0.00
bn140320092 2.30 5.60 281.86 -11.17 0.242 17.94 281.97 -6.71 902.86 0.00 0.00
2014-03-23 10:24:39.39 - - - - 1.086 10.72 11.30 -75.91 110.41 0.00 0.00
2014-03-23 20:48:42.70 - - - - 2.671 10.08 313.85 53.01 60.57 0.00 0.00
bn140329272 0.06 10.70 92.35 -41.08 0.073 32.42 94.57 -46.41 366.86 0.00 0.00
bn140402007 0.32 5.70 207.59 5.97 0.098 20.52 200.90 7.90 902.86 0.00 0.00
2014-04-05 14:44:14.78 - - - - 1.979 9.81 326.22 -12.30 271.72 0.00 0.03
2014-04-12 17:31:24.51 - - - - 0.327 10.93 243.31 -28.21 902.86 0.00 0.00
bn140428906 0.32 7.40 2.01 68.17 0.133 37.84 358.58 62.82 668.72 0.00 0.00
bn140501139 0.26 - 171.88 24.64 0.133 15.49 167.47 30.51 495.30 0.00 0.00
bn140511095 1.41 10.70 329.76 -30.06 0.098 23.57 313.76 -25.40 271.72 0.00 0.00
2014-05-13 03:18:50.18 - - - - 3.606 9.06 178.49 66.61 110.41 0.06 0.00
bn140518709 0.70 4.70 244.04 -77.86 0.133 11.21 277.42 -64.21 366.86 0.00 0.00
bn140526571 0.06 6.00 142.99 -10.95 0.098 17.60 145.65 -15.57 668.72 0.00 0.00
bn140605377 0.51 8.30 121.79 -53.86 0.179 47.37 122.45 -60.64 668.72 0.00 0.00
2014-06-06 10:58:13.56 - - - - 0.327 12.98 224.26 38.46 201.26 0.00 0.00
bn140610487 0.96 6.20 199.05 35.91 0.179 16.51 199.57 35.10 366.86 0.00 0.00
2014-06-10 $1$ 1.086 13.27 219.84 -25.70 44.86 0.00 0.00
bn140619475 2.82 5.60 132.70 -9.70 0.596 42.96 135.77 -4.31 902.86 0.00 0.00
bn140619490 0.45 5.80 304.96 -39.18 0.133 17.83 247.45 -23.79 271.72 0.00 0.00
2014-06-22 11:11:55.76 - - - - 0.073 9.54 281.88 -9.43 44.86 0.54 0.00
bn140624423 0.10 7.70 23.16 -0.56 0.022 58.16 22.39 -10.07 271.72 0.00 0.00
bn140626843 1.79 6.10 120.81 38.78 1.466 31.69 125.61 42.09 110.41 0.00 0.00
2014-07-01 18:38:26.69 - - - - 0.098 11.41 38.08 -21.99 110.41 0.00 0.00
2014-07-04 19:24:42.31 - - - - 1.466 8.63 301.99 -33.86 110.41 0.00 0.00
2014-07-07 05:59:44.82 - - - - 3.606 9.09 320.46 9.21 149.06 0.00 0.00
bn140710537 0.38 5.60 2.80 -38.88 0.441 17.12 357.43 -31.08 149.06 0.34 0.00
bn140716306 3.33 5.20 215.23 57.01 2.671 19.76 231.55 47.12 201.26 0.00 0.00
2014-07-16 10:29:38.39 - - - - 1.086 16.77 106.97 -56.84 110.41 0.00 0.00
2014-07-19 05:28:18.47 - - - - 1.086 9.36 35.92 55.97 60.57 0.82 0.00
bn140720158 0.32 7.80 175.03 -32.31 0.022 17.81 202.31 -20.95 201.26 0.00 0.00
bn140724533 0.90 5.20 314.73 -1.85 0.441 11.30 299.43 -8.44 110.41 0.00 0.00
2014-08-01 18:58:09.43 - - - - 1.979 8.82 19.34 15.66 44.86 0.07 0.00
bn140807500 0.51 5.20 200.16 26.49 0.327 66.51 195.93 32.20 495.30 0.00 0.00
2014-08-25 07:53:51.22 - - - - 0.441 9.68 5.06 23.70 201.26 0.00 0.00
bn140831215 0.70 4.70 280.37 25.64 0.242 10.59 278.58 38.44 271.72 0.00 0.00
bn140831374 3.58 4.80 4.29 44.01 1.086 28.84 351.03 39.80 366.86 0.00 0.00
bn140901821 0.18 13.40 15.82 -32.76 0.133 206.81 16.11 -32.36 1645.78 0.00 0.00
bn140930134 3.26 4.70 41.61 57.58 1.466 20.49 21.03 33.66 366.86 0.03 0.00
bn141004973 2.56 4.80 76.75 12.82 0.441 33.85 62.52 15.80 81.78 0.00 0.00
bn141005217 3.39 5.00 291.09 36.10 2.671 53.42 292.81 44.17 110.41 0.00 0.00
bn141011282 0.08 35.00 257.94 -9.68 0.179 64.23 251.51 1.28 495.30 0.00 0.00
bn141020439 1.60 4.60 265.49 -1.68 0.804 10.98 268.31 -11.14 495.30 0.00 0.00
2014-10-23 04:24:57.77 - - - - 0.179 11.28 195.86 -68.46 60.57 0.00 0.00
bn141026742 2.56 5.70 132.82 62.48 1.979 22.03 102.92 53.45 201.26 0.00 0.00
2014-10-27 03:24:33.67 - - - - 0.073 9.58 93.03 -59.05 44.86 0.00 0.00
bn141031998 0.16 6.20 133.08 -33.68 0.133 13.79 117.16 -29.82 366.86 0.02 0.00
bn141102112 0.02 5.10 223.23 -17.42 0.030 15.25 220.87 3.27 201.26 0.04 0.53
bn141102536 2.62 7.80 208.61 -47.10 0.441 43.31 212.95 -50.83 366.86 0.00 0.00
bn141105406 1.28 6.40 16.91 29.17 0.596 57.53 18.44 28.67 495.30 0.00 0.00
bn141111435 1.73 4.80 51.22 43.02 0.327 10.15 34.53 56.21 668.72 0.00 0.00
bn141113346 0.45 5.20 171.02 80.26 0.441 16.63 200.65 67.39 902.86 0.01 0.00
2014-11-20 11:29:29.44 - - - - 3.606 10.64 150.31 -49.43 44.86 0.15 0.00
bn141122087 1.28 5.70 9.71 -20.02 0.098 17.08 8.01 -5.38 495.30 0.00 0.00
bn141124277 0.51 4.90 135.07 78.18 0.441 17.49 145.34 76.03 495.30 0.00 0.00
bn141126233 0.90 4.80 243.87 59.99 0.596 11.66 247.34 61.45 271.72 0.00 0.00
bn141128962 0.27 12.70 321.80 -35.76 0.054 30.90 317.84 -34.01 110.41 0.00 0.00
bn141202470 1.41 5.60 143.07 54.16 1.086 79.11 132.93 54.76 668.72 0.00 0.00
bn141205337 1.28 5.10 92.86 37.88 0.596 20.64 82.10 -5.26 495.30 0.00 0.00
bn141208632 0.96 5.90 359.26 26.44 0.073 13.17 356.40 38.08 366.86 0.00 0.00
bn141213300 0.77 5.90 248.19 18.06 0.242 24.91 254.61 23.37 110.41 0.00 0.00
2014-12-17 04:22:13.27 - - - - 0.804 9.45 95.72 32.23 495.30 0.03 0.00
2014-12-19 06:47:56.80 - - - - 2.671 8.88 196.55 -2.85 110.41 0.00 0.00
bn141222298 2.75 9.90 178.04 -57.35 3.606 158.69 169.17 -57.20 495.30 0.00 0.00
2014-12-26 01:00:08.93 - - - - 2.671 12.50 75.02 -18.39 81.78 0.02 0.00
bn141230871 0.22 6.70 246.93 -40.18 0.098 17.69 240.60 -46.35 33.23 0.00 0.00
bn140109877 3.33 5.10 24.09 -25.05 - - - - - - -
bn140616165 0.51 5.60 104.94 -70.51 - - - - - - -
bn140912664 2.30 - 303.50 59.56 - - - - - - -

### v2

@@ Appendix I Contaminant Removal Tests Histograms
NameaaName: The trigger name as appears in the GBM trigger catalog. If it is a new trigger, the trigger time is displayed instead. T90 (s) SNR ($\sigma$)bbSNR ($\sigma$): The significance of the trigger at triggering, as appears in the GBM data product: TRIGDAT file. It is reported in units of standard deviations of standard normal distribution, hence it has the same meaning as our SNR. RA (deg) Dec (deg) Duration (s)ccDuration: The burst duration that had the maximal SNR in our search. SNR ($\sigma$) RA (deg)ddRA (deg): The maximum likelihood estimation of the right ascension, obtained by the sky posterior. Dec (deg)eeDEC (deg): The maximum likelihood estimation of the declination, obtained by the sky posterior. $E_{\text{peak}}$ (KeV)ffSubthreshold/Other: Indicated if another subthreshold search or other instrument has detected this candidate. $\mathcal{B}_{\oplus}$ $\mathcal{B}_{\odot}$
2014-01-01 07:50:54.12 - - - - 2.671 8.42 325.31 22.44 44.86 0.00 0.00 -
bn140105065 1.09 6.40 208.22 50.17 0.242 21.44 210.17 49.79 366.86 0.00 0.00 -
bn140105748 0.58 4.70 252.88 19.03 3.606 19.15 256.83 23.13 495.30 0.00 0.00 -
bn140109771 0.70 4.80 102.74 29.76 0.327 16.21 195.73 41.64 668.72 0.00 0.00 -
bn140110411 0.77 4.60 50.64 -69.29 0.327 11.06 334.96 -76.88 366.86 0.30 0.20 -
bn140122597 3.58 4.80 56.08 15.09 2.671 17.58 58.98 3.16 110.41 0.00 0.00 -
2014-01-24 18:00:09.62 - - - - 0.133 11.90 222.52 -36.77 495.30 0.00 0.00 GBM subthreshold
2014-01-29 07:06:50.56 - - - - 0.133 10.79 79.91 -6.35 201.26 0.00 0.00 -
bn140129499 0.13 8.20 183.40 -10.32 0.054 23.71 185.10 -3.97 149.06 0.00 0.00 -
2014-02-02 21:48:46.53 - - - - 0.022 8.39 272.66 88.03 110.41 0.98 0.00 -
bn140209313 1.41 5.40 81.33 32.49 0.596 235.61 89.66 23.69 110.41 0.00 0.00 -
2014-02-09 22:03:12.19 - - - - 2.671 9.78 281.68 -28.26 81.78 0.00 0.07 -
bn140211091 3.46 5.00 115.84 -13.59 1.979 18.46 115.91 -10.83 110.41 0.00 0.00 -
bn140216331 2.43 5.10 194.04 31.46 0.596 12.25 204.93 39.76 366.86 0.00 0.00 -
bn140224382 2.30 4.70 2.78 20.37 2.671 25.91 330.99 27.19 271.72 0.00 0.00 -
2014-02-25 04:03:20.64 - - - - 3.606 10.03 348.73 17.04 149.06 0.01 0.00 -
2014-03-03 06:44:35.49 - - - - 0.073 12.02 297.48 8.78 81.78 0.00 0.00 -
2014-03-07 00:43:25.41 - - - - 0.022 17.68 148.46 69.51 1645.78 0.09 0.00 GBM subthreshold
bn140320092 2.30 5.60 281.86 -11.17 0.242 17.94 281.97 -6.71 902.86 0.00 0.00 -
2014-03-23 20:48:42.70 - - - - 2.671 10.08 313.85 53.01 60.57 0.00 0.00 -
bn140329272 0.06 10.70 92.35 -41.08 0.073 32.42 94.57 -46.41 366.86 0.00 0.00 -
bn140402007 0.32 5.70 207.59 5.97 0.098 20.52 200.90 7.90 902.86 0.00 0.00 -
2014-04-05 14:44:14.78 - - - - 1.979 9.81 326.22 -12.30 271.72 0.00 0.03 -
2014-04-12 17:31:24.51 - - - - 0.327 10.93 243.31 -28.21 902.86 0.00 0.00 GBM subthreshold
bn140428906 0.32 7.40 2.01 68.17 0.133 37.84 358.58 62.82 668.72 0.00 0.00 -
bn140501139 0.26 - 171.88 24.64 0.133 15.49 167.47 30.51 495.30 0.00 0.00 -
bn140511095 1.41 10.70 329.76 -30.06 0.098 23.57 313.76 -25.40 271.72 0.00 0.00 -
2014-05-13 03:18:50.18 - - - - 3.606 9.06 178.49 66.61 110.41 0.06 0.00 -
bn140518709 0.70 4.70 244.04 -77.86 0.133 11.21 277.42 -64.21 366.86 0.00 0.00 -
bn140526571 0.06 6.00 142.99 -10.95 0.098 17.60 145.65 -15.57 668.72 0.00 0.00 -
2014-05-27 14:56:06.45 - - - - 1.466 8.45 357.72 0.40 81.78 0.02 0.00 -
bn140605377 0.51 8.30 121.79 -53.86 0.179 47.37 122.45 -60.64 668.72 0.00 0.00 -
2014-06-06 10:58:13.56 - - - - 0.327 12.98 224.26 38.46 201.26 0.00 0.00 GBM subthreshold, Swift, ACS
bn140610487 0.96 6.20 199.05 35.91 0.179 16.51 199.57 35.10 366.86 0.00 0.00 -
2014-06-10 16:18:39.11 - - - - 1.086 13.27 219.84 -25.70 44.86 0.00 0.00 GBM subthreshold
bn140619475 2.82 5.60 132.70 -9.70 0.596 42.96 135.77 -4.31 902.86 0.00 0.00 -
bn140619490 0.45 5.80 304.96 -39.18 0.133 17.83 247.45 -23.79 271.72 0.00 0.00 -
2014-06-22 11:11:55.76 - - - - 0.073 9.54 281.88 -9.43 44.86 0.54 0.00 -
bn140624423 0.10 7.70 23.16 -0.56 0.022 58.16 22.39 -10.07 271.72 0.00 0.00 -
bn140626843 1.79 6.10 120.81 38.78 1.466 31.69 125.61 42.09 110.41 0.00 0.00 -
2014-07-01 18:38:26.69 - - - - 0.098 11.41 38.08 -21.99 110.41 0.00 0.00 GBM subthreshold
2014-07-04 19:24:42.31 - - - - 1.466 8.63 301.99 -33.86 110.41 0.00 0.00 -
2014-07-07 05:59:44.82 - - - - 3.606 9.09 320.46 9.21 149.06 0.00 0.00 -
bn140710537 0.38 5.60 2.80 -38.88 0.441 17.12 357.43 -31.08 149.06 0.34 0.00 -
bn140716306 3.33 5.20 215.23 57.01 2.671 19.76 231.55 47.12 201.26 0.00 0.00 -
2014-07-19 05:28:18.47 - - - - 1.086 9.36 35.92 55.97 60.57 0.82 0.00 -
bn140720158 0.32 7.80 175.03 -32.31 0.022 17.81 202.31 -20.95 201.26 0.00 0.00 -
bn140724533 0.90 5.20 314.73 -1.85 0.441 11.30 299.43 -8.44 110.41 0.00 0.00 -
bn140807500 0.51 5.20 200.16 26.49 0.327 66.51 195.93 32.20 495.30 0.00 0.00 -
bn140831215 0.70 4.70 280.37 25.64 0.242 10.59 278.58 38.44 271.72 0.00 0.00 -
bn140831374 3.58 4.80 4.29 44.01 1.086 28.84 351.03 39.80 366.86 0.00 0.00 -
bn140901821 0.18 13.40 15.82 -32.76 0.133 206.81 16.11 -32.36 1645.78 0.00 0.00 -
bn140930134 3.26 4.70 41.61 57.58 1.466 20.49 21.03 33.66 366.86 0.03 0.00 -
bn141004973 2.56 4.80 76.75 12.82 0.441 33.85 62.52 15.80 81.78 0.00 0.00 -
bn141005217 3.39 5.00 291.09 36.10 2.671 53.42 292.81 44.17 110.41 0.00 0.00 -
bn141011282 0.08 35.00 257.94 -9.68 0.179 64.23 251.51 1.28 495.30 0.00 0.00 -
bn141020439 1.60 4.60 265.49 -1.68 0.804 10.98 268.31 -11.14 495.30 0.00 0.00 -
2014-10-21 14:54:13.07 - - - - 0.016 8.46 348.43 14.37 60.57 0.00 0.00 -
2014-10-23 04:24:57.77 - - - - 0.179 11.28 195.86 -68.46 60.57 0.00 0.00 -
bn141026742 2.56 5.70 132.82 62.48 1.979 22.03 102.92 53.45 201.26 0.00 0.00 -
2014-10-27 03:24:33.67 - - - - 0.073 9.58 93.03 -59.05 44.86 0.00 0.00 -
bn141031998 0.16 6.20 133.08 -33.68 0.133 13.79 117.16 -29.82 366.86 0.02 0.00 -
bn141102112 0.02 5.10 223.23 -17.42 0.030 15.25 220.87 3.27 201.26 0.04 0.53 -
bn141102536 2.62 7.80 208.61 -47.10 0.441 43.31 212.95 -50.83 366.86 0.00 0.00 -
bn141105406 1.28 6.40 16.91 29.17 0.596 57.53 18.44 28.67 495.30 0.00 0.00 -
bn141111435 1.73 4.80 51.22 43.02 0.327 10.15 34.53 56.21 668.72 0.00 0.00 -
bn141113346 0.45 5.20 171.02 80.26 0.441 16.63 200.65 67.39 902.86 0.01 0.00 -
2014-11-20 11:29:29.44 - - - - 3.606 10.64 150.31 -49.43 44.86 0.15 0.00 GBM subthreshold
bn141122087 1.28 5.70 9.71 -20.02 0.098 17.08 8.01 -5.38 495.30 0.00 0.00 -
bn141124277 0.51 4.90 135.07 78.18 0.441 17.49 145.34 76.03 495.30 0.00 0.00 -
bn141126233 0.90 4.80 243.87 59.99 0.596 11.66 247.34 61.45 271.72 0.00 0.00 -
bn141128962 0.27 12.70 321.80 -35.76 0.054 30.90 317.84 -34.01 110.41 0.00 0.00 -
bn141202470 1.41 5.60 143.07 54.16 1.086 79.11 132.93 54.76 668.72 0.00 0.00 -
bn141205337 1.28 5.10 92.86 37.88 0.596 20.64 82.10 -5.26 495.30 0.00 0.00 -
bn141208632 0.96 5.90 359.26 26.44 0.073 13.17 356.40 38.08 366.86 0.00 0.00 -
bn141213300 0.77 5.90 248.19 18.06 0.242 24.91 254.61 23.37 110.41 0.00 0.00 -
2014-12-17 04:22:13.27 - - - - 0.804 9.45 95.72 32.23 495.30 0.03 0.00 -
2014-12-19 06:47:56.80 - - - - 2.671 8.88 196.55 -2.85 110.41 0.00 0.00 -
bn141222298 2.75 9.90 178.04 -57.35 3.606 158.69 169.17 -57.20 495.30 0.00 0.00 -
2014-12-26 01:00:08.93 - - - - 2.671 12.50 75.02 -18.39 81.78 0.02 0.00 GBM subthreshold
bn140109877 3.33 5.10 24.09 -25.05 - - - - - - - -
bn140616165 0.51 5.60 104.94 -70.51 - - - - - - - -
bn140912664 2.30 - 303.50 59.56 - - - - - - - -
bn141230871 0.22 6.70 246.93 -40.18 - - - - - - - -